HUBBLE SPACE TELESCOPE DAILY REPORT #833 PERIOD COVERED: 0800 EST Daily Status Report as of 060/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of PC 4714 (WF/PC Calibration: Internal Delta Flats) The Planetary Camera was used to make internal flat observations using filter F673N. This image will be used in the restoration of Earth flats, which are altered due to changes in quantum efficiency which occurred after decontamination operations. The observations were executed as scheduled, and monitored in the OSS. No problems were reported. 1.2 Completed GHRS/2 3584 (Probing the Voids II) The High Resolution Spectrograph was used to make observation of an extragalactic void by studying the background active galactic nucleus of galaxy Markarian 501. The observations were reported to be weak and noisy. However, the Principal Investigator said that this was expected. 1.3 Completed Two Sets of WFC 4713 (WF/PC Calibration: Internal Delta Flats) The Wide Field Camera was used to make a total of four internal flat observations. These observations will be used for the restoration of Earth flats, which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. All four images were reported to look okay. 1.4 Completed HSP/POL 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make two ultraviolet observations of standard star VID998 to measure the sensitivity of its detector tubes. The observations were executed as scheduled, and no problems were reported. The images were well centered, and reported to look okay. 1.5 Completed HSP/UV2 1095 (Variability of High Luminosity Stars) The High Speed Photometer was used to make three observations of one of the most luminous and massive stars known, HD93250. The observations were monitored in the OSS, and no problems were reported. 1.6 Completed FOC/96 3591 (Massive Stars in Starburst Galaxies- FOC Images Part 1) The Faint Object Camera (f/96) was used to make observations of the star burst regions in galaxy NGC1705 to study the process of the galaxy's formation and evolution. The observations were monitored in the OSS, and no problems were reported. 1.7 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the wide Field Camera and the Planetary Camera were used to make internal flat observations to look for new forms of contamination (i.e., measles and daisies), and the return of quantum efficiency hysteresis. All of the observations were reported to look normal. 1.8 Completed Two Sets of FOS/RD 3744 (Spectrophotometry of Phobos and Deimos) The Faint Object Spectrograph was used to make two sets of Mars' moon Deimos. The first observations was reported to look normal. On the second observation the OSS attempted to uplink a pointing correction to eliminate the effects of changes in the HST orbital ephemeris. The OSS was unable to uplink these commands (OSIEVENT(SIP) from SCIF), and they were sent from the MOR (POCC). Subsequent observations were reported to look okay. 1.9 Completed Three Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make observations of the protoplanetary system PK 358-00D2 to study its current morphology. All three observations were reported to look normal. 1.10 Completed PC 4347 (WF/PC Calibration: Flood Retention Test) The Planetary Camera was used to make an Earth flat image using filter F439W to determine how much of the last ultraviolet flood still remains in the system. The observations were reported to look normal. 1.11 Completed PC 4093 (Survey of Nuclear Regions of 102 Markarian Galaxies) The Planetary Camera was used to make one image of the inner regions of Markarian Galaxy UGC4079. The observation was executed in course track. The image was received and displayed in real-time. It was reported to look okay. 1.14 Uplink Updated Ephemeris Tables, Per MOSES 2871 Uplinked at 058/0151Z and 060/0023Z. 1.15 Complete FOS/RD 4220 (Imaging and Spectrophotometry of Seyfert Nuclei) The Faint Object Spectrograph was used to take four spectral observations of Seyfert Galaxy NGC-1068. All of the observations were made in course track. The target was reported to be well centered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 13 Successful : 13 Scheduled Reacquisitions : 32 Successful : 32 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4123 (FOS Spectral Flat Fields Monitor Revised) 3.2 Complete PC 4712 (WF/PC Calibration: Measles Monitor) 3.3 Complete WFC/FOS/BL 3707 (HST Observations of Periodic Comets) 3.4 Complete PC 4093 (Survey of Nuclear Regions of 102 Markarian Galaxies) 3.5 Complete PC 4027 (Non-Proprietary Snapshots Survey Round 2) 3.6 Complete WFC 3172 (WF/PC SV Observations: Internal Monitor Proposal) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-